NBIPComputing's profile picture. Supporting world-class science at the Norwich BioScience Institutes

NBIP Computing

@NBIPComputing

Supporting world-class science at the Norwich BioScience Institutes

Power has been restored to the core of the network in JIC B30. Services restored.


A power problem has taken out the core network in JIC B30. We hope to have power restored very soon.


Networking in B30 working normally again...


A network device has failed on the 1st floor of B30. This is affecting some wireless access points and ~10 desktop ports. We are prepping a replacement; all will be operational again shortly.


As part of planned improvements to the NBI network, essential maintenance needs to be carried out on Saturday 30 June 2018, between 08:00 – 10:00. Disruption will be kept to a minimum with only brief intermittent loss to internet access expected during this time.


Issue with JIC telephone line at the moment. Outgoing routing via IFR. Incoming not working.


There seems to be an issue with the JANET network causing certain sites to be less responsive than normal. They are investigating.


SMTP issue resolved 10am. Issue due to server changes at Microsoft.


We're investigating an issue with sending SMTP (automated) mails. Standard emails are not affected.


Apple have released security updates for both macOS Sierra (10.12.3) and iOS (10.2.1),addressing multiple *critical* vulnerabilities.


for our twitter followers: a sneak peak of the new intranet - go to http://intranet2 going live early 2017


Having problems with an iPhone connected to eduroam? Might be iOS10. See http://intranet/infoserv/cgi-bin/nb/detail.asp?id=27664&Board=3


Eduroam issues resolved in. Conf. Centre. Unfortunately fix is disruptive so having to be careful about applying to the rest of the site.


We're investigating an issue with Eduroam affecting some areas of the site


Telephony fault confirmed to be at the exchange. BT will fix either this PM or tomorrow AM.


Outgoing calls are sharing IFR's trunk, so will work as long as they are not overwhelmed at the time.


We are investigating an issue with the telephone system. Looks like the EI/JIC/TSL BT trunk might be down. This means no incoming calls.


Flood in B26 server room yesterday. Primary isilon and all cluster services affected. CiS will give restoration ETA when known.


SQL server working as normal now.


our main SQL server is unhappy about something. Affecting intranet & other services. We're investigating...


United States Trends

Loading...

Something went wrong.


Something went wrong.